THE FORBIDDEN ROTATION AND ROTATION-VIBRATION SPECTRUM OF H_{2}^{+}

\underline{P. R. BUNKER}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada; R. E. MOSS, Department of Chemistry, University of Southampton, Southampton SO17 1BJ, U.K..

Previous studies of the g/u electronic symmetry breaking interaction between the rotation-vibration levels of the X^{2}Σ_{g}^{+} ground electronic state and the A^{2}Σ_{u}^{+} first excited electronic state of H_{2}^{+} have focused on explaining the large hyperfine structure observed in some rovibronic transitions between these electronic states^{a}^{,}^{b}. The interaction is an ortho-para interaction between levels of the two electronic states. It gives rise to an interesting new electric dipole spectrum that consists of forbidden rotation and rotation-vibration transitions within the X^{2}Σ_{g}^{+} ground state involving levels that are near the dissociation limit^{c}. Such a spectrum will occur for any homonuclear diatomic ion having nuclei with non-zero spin such as ^{3}He_{2}^{+}, Li_{2}^{+} and ^{21}Ne_{2}^{+}.

aR. E. Moss, Chem. Phys. Lett. \underline{206}, 83 (1993), and references therein.

bSection 17.7 of the book Molecular Symmetry and Spectroscopy, 2nd Edition, by P. R. Bunker and P. Jensen, NRC Research Press, Ottawa 1998. See http://www.nrc.ca/cisti/journals/41653 for the table of contents and ordering information.

cP. R. Bunker and R. E. Moss, Chem. Phys. Lett. \underline{316}, 266 (2000).